which of the following is the greatest common factor of the terms 48s^5t^3u^ and 40s^3tu^4?
Answer by stanbon(75887)

which of the following is the greatest common factor of the terms 48s^5t^3u^ and 40s^3tu^4?
-----------------------------------
gcf of 48 and 40 = 8
gcf of s^5 and s^3 = s^3
gcf of t^3 and t = t
gcf of u^4 and u^? = unknown
=================================
GCF = 8s^3tu^(unknown)
=============================
Cheers,
Stan H.